OpenStudy (anonymous):

What is the rate per interest period of 12% compounded quarterly?

OpenStudy (welshfella):

how many quarters are there in 1 year?

OpenStudy (anonymous):

Four.

OpenStudy (anonymous):

@Nalemi P is the principal (the initial amount you borrow or deposit) r is the annual rate of interest (percentage) n is the number of years the amount is deposited or borrowed for. A is the amount of money accumulated after n years, including interest. When the interest is compounded once a year: A = P(1 + r)n

OpenStudy (anonymous):

rate per interest period shall be 12/4 = 3% per quarter

OpenStudy (anonymous):

rate per interest period shall be 12/4 = 3% per quarter